OpenStudy (anonymous):

help factoring x^6-x^12

OpenStudy (owlcoffee):

There are several ways to factor this, but, let's multipy and divide by x^6 in order to make things more simple, this woud be a variant of taking common factor: \[x^6-x ^{12} \iff \frac{ x^6 }{ x^6 }(x^6-x ^{12}) \iff x^6(\frac{ x^6 }{ x^6 }-\frac{ x ^{12} }{ x^6 })\] \[\iff x^6(1-x^6)\] Now we have to acomodate it so it becomes a difference of squares: \[x^6(1-x^6) \iff x^6(1-(x^3)^2)\] The rest is pretty straight forward, you just have to apply the structure: \(a^2-b^2=(a+b)(a-b)\) and remember that \(1^2=1\).

OpenStudy (anonymous):

well... the answer is supposed to be (x-y^2)(x^2+xy^2+y^4)(x+y^2)(x^2-xy^2+y^4) :/

OpenStudy (owlcoffee):

Make sure you properly type the exercise next time.